OpenStudy (anonymous):

Buses arrive at a resort hotel every 30mins. They wait for 5 mins while passengers get off and on, and then the buses depart. what is the probability that there is a bus waiting when a hotel guest walks out of the door at a randomly chosen time?

OpenStudy (anonymous):

Bus waiting time divided by the total time. total time = 30 +5 = 35. So 5/35= 1/7= about 14.3%

OpenStudy (anonymous):

can you help me with some of the questions i posted?? please

OpenStudy (anonymous):

Ok. Post here.

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

Find the probability that a randomly chosen point in the figure lies in the shaded region...

OpenStudy (anonymous):

theres this one and one more

OpenStudy (anonymous):

Find the probability that a point X, selected randomly on line LQ, is on the given segment. 1)Line LM 2)line NP 3)line OQ 4)line MQ... fractions simplify

OpenStudy (anonymous):

line LM is 2/24 = 1/12 = 8.33% line NP is 12/24 = 1/2 = 50.0% line OQ is 12/24 = 50.0% line MQ is 22/24 = 91.67%

OpenStudy (anonymous):

wouldnt it be 2/17?

OpenStudy (anonymous):

Picture 21: area of a circle = Pi *radius * radius. The radius of a circle is a half of a half or 13/4 =3.25. because you have 4 circles it is 4 * Pi * radius * radius = about 132.732. the total area of the square is 13*13 = 169. so the white part is 132.732/169 = 78.54%. and the black 36.27/169 = 21.46%.

OpenStudy (anonymous):

line Lm wouldn't be 2/17 because you need to consider the other number to the left... or the negative numbers. because each line are two units apart and there are 12 of them the total is then 24.

OpenStudy (anonymous):

wow, you guys are greate